0 0
Add a comment Improve this question Transcribed image text
Answer #1

apply I/io = e^-t/RC

i/i0 = (e^-(0.01/(369*0.215e-6)

i/i0 = 1.811 Amps

--------------------------------------------

apply impedence Z^2 = ( R^2+XL^2)

so R^2 = 55.1^2 -(26^2)

R^2 = 2360.61

R = 48.57 ohms